ASTROCHEMISTRY WITH THE COMBINED ARRAY FOR RESEARCH IN MILLIMETER-WAVE AS-TRONOMY

\underline{D. N. FRIEDEL}, L. W. LOONEY, Department of Astronomy, University of Illinois, Urbana, IL 61801.

The Combined Array for Research in Millimeter-wave Astronomy (CARMA) is now actively observing at λ=3mm frequencies (85-115 GHz) and commissioning of the λ=1mm (210-270 GHz) system is ongoing. The spectroscopic and imaging capabilities of this state of the art instrument will be presented. We acknowledge support from the Laboratory for Astronomical Imaging at the University of Illinois and NSF grant AST-0540459.